# Chameleon Channels: Measuring YouTube Accounts Repurposed for Deception and Profit

### v1

@@ 3.1 The Fameswap Market
Fameswap is an online marketplace for buying and selling accounts on YouTube, TikTok, Instagram, Twitter/X, and websites. Recent measurements identify it as the largest such market by number of sellers [7]. The site resembles a standard e-commerce platform: each account has a dedicated listing with a description, price, seller reviews, and a content category (e.g., humor, sports). Buyers can bid, and the platform offers escrow, dispute resolution, and account verification—features similar to marketplaces like eBay.
Fameswap earns revenue from escrow fees (3% or a USD 50 minimum) and premium accounts, which offer advanced search, additional metrics, lower fees, and more. Payments are accepted via PayPal, wire transfers, and cryptocurrencies (e.g., BTC, ETH, USDC).
@@
As of writing, 25,404 listings across all platforms are advertised for a combined total of USD 366.7M, claiming over 2.6B followers. For YouTube, we filtered for unique channel IDs, identifying 4,641 listings with 823M subscribers (confirmed via the YouTube API by March 31, 2025) and a total listed price of USD 160.4M. Chu et al. found 3,112 listings between 2019–2021, suggesting a nearly 9x growth [8]. The average listing price is USD 5,400 with 105,751 subscribers—comparable to Chu et al.’s findings. Like Beluri et al., we observed inflated pricing: 41 listings exceeded USD 100K and 16 exceeded USD 1M.
@@
To contextualize these prices, we computed the cost per 1,000 subscribers—a common metric on engagement forums [18]. As shown in Figure 3, this varies widely by category. The median cost on Fameswap is substantially higher than the USD 16.52 per 1,000 subscribers found by Nevado-Catalán et al. [18]

### v2

@@ 4.4 Scaling Annotations with LLMs
Content Labels. Similarly, using a prompt based on the annotation codebook we developed, we annotated three samples: repurposed Fameswap channels, repurposed Social Blade channels, and a random sample of non-repurposed Social Blade channels which we will use as a control group. For each channel, we provide the LLM with up to 50K tokens per channel, combining the channel description with video titles and descriptions. Most channels fit this limit. If a channel exceeds it, we (i) de-duplicate strings with >> 0.9 edit similarity to capture more diverse content, then (ii) randomly sample the remaining text until the total is $\leq$50K tokens. We do this to capture a greater diversity of information, given that some channels post many videos with the same or similar titles. We cap input at 50K (well below GPT-4o’s 128 K window) because longer contexts seem to degrade classification accuracy [51].
For each of these samples, we annotate the content of a random 5% (nn=260, out of nn=5,098) subsample drawn from the three samples. These channels had a combined total of 14,354 videos, which the lead author manually annotated. The annotations to the content are binary, indicating the presence of content.
5 Fameswap Results
@@
We found 25,404 listings for social media accounts across all platforms, advertised worth a total of USD 366.7M and claiming over 2.6B followers. Between 2019-2021, Chu et al. found 3,112 listings on Fameswap [15], suggesting substantial growth to the present day. For YouTube specifically, we found 4,641 listings with 823M subscribers (confirmed via the YouTube API) and a combined listed price of USD 160.4M.
@@
To contextualize listing prices, we computed the cost per 1,000 subscribers—a common metric in engagement forums [60]. As shown in Figure 4, this varies widely by category. The median cost on Fameswap in some categories is substantially higher than USD 16.52 per 1,000 subscribers found by Nevado-Catalán et al. [60], suggesting that other features like niche drive the valuation and that the audience for sale is pricier than a comparable sized inorganic audience.